Can you help me solve this question #Assignment# A spring of Force constant 1500Nm-¹ is acted upon by a constant force of 45N . calculate the potential energy stored in spring

1
Expert's answer
2022-02-09T10:50:13-0500

Given:

"k=\rm 1500\:\rm N/m"

"F=45\:\rm N"


The deformation of spring

"x=\frac{F}{k}=\frac{45\:\rm N}{1500\:\rm N/m}=0.03\:\rm m"

The potential energy stored in spring

"E_p=\frac{kx^2}{2}=\frac{1500*0.03^2}{2}=0.675\:\rm J"